A 1.25-g sample of CO2 is contained in a 750.-mL flask at 22.5 °C. What is the pressure

of the gas?


1
Expert's answer
2021-10-16T02:01:52-0400

"m=1.25 \;g \\

V=750 \;mL = 0.75 \;L \\

T= 22.5+273.15 = 295.65 \;K \\

M(CO_2) = 44.01 \;g/mol \\

n(CO_2) = \frac{1.25}{44.01} = 0.0284 \;mol"

Ideal gas law

pV = nRT

R = 0.08206 L×atm/mol×K

"p=\frac{nRT}{V} \\

p= \frac{0.0284 \times 0.08206 \times 295.65}{0.75} = 0.918 \;atm"

Answer: 0.918 atm